Find an equation of the line tangent to the curve y=3x²-1 and perpendicular to the line x-3y=4


1
Expert's answer
2021-10-15T04:56:22-0400

Let y=g(x) be a sought line.

The equation of y=g(x) could be found the following way: "g(x) - g{\scriptscriptstyle 0} = k{\scriptscriptstyle 1}(x-x{\scriptscriptstyle 0})", where ("x{\scriptscriptstyle 0},y{\scriptscriptstyle 0})" is the point belonging to a line, "k{\scriptscriptstyle 1}" is the slope of the line.

Since y=g(x) is perpendicular to x-3y=4, then "k{\scriptscriptstyle 1}k{\scriptscriptstyle 2}=-1", where "k{\scriptscriptstyle 2}" is the slope of the second line.

"x-3y=4\to 3y=x-4\to y = {\frac x 3} - {\frac 4 3} \to k{\scriptscriptstyle 2} = {\frac 1 3}"

"{\frac 1 3}k{\scriptscriptstyle 1} = -1 \to k{\scriptscriptstyle 1} = -3"

g(x) is tangent to "y = 3x^{2} -1". The slope for the tangent line to "y = 3x^{2} -1" is "y'=6x"

Then "6x_0 = -3"

"x{\scriptscriptstyle 0}" = -0.5

"y{\scriptscriptstyle 0} = 3*(-0.5)^{2} -1 = -0.25"

Then "g(x) + 0.25 = -3(x +0.5) \to g(x) = -3x - 1.75"

So, the equation of a sought line is "4y + 12x + 7" = 0